# Empirical bounds for commuting dilations of free unitaries and the universal commuting dilation constant

### v1

@@ 1. Introduction
In recent years, dilation theoretic techniques and the framework of matrix convex sets have found new and significant applications to quantum information [BN18, BN20, BNS23], optimization and control [HKMS19], mathematical physics [ALPP21, GS23], the theory of operator systems and operator algebras [DP20, FNT17, PP21], and more [DlCDN20, DlCNVB23, EPŠ25, HKM17]. The goal of this paper is to advance our knowledge regarding an open problem in this area: determining of the universal commuting dilation constant $C_{2}$, i.e. the smallest number $c$ such that every pair of contractions is the simultaneous compression of commuting normal operators of norm at most $c$. By combining rigorous limit theorems with numerical experimentation, we provide evidence that $C_{2}\leq 2\sqrt{\frac{2}{3}}<2$.
@@ 1.1. Preliminaries and background
The one sided matrix range distance $\operatorname{d}_{\mathrm{mr}}(B\to A)$ was discussed before in [DDSS17, Section 5] where it was denoted by $\delta_{\mathcal{W}(A)}(\mathcal{W}(B))$. The matrix range distance is then given by
@@
The matrix range distance is a metric only when considered on classes of certain rigid tuples, for example on the set $\mathcal{U}(d)$ of unitary $d$-tuples. For general operator tuples it does not determine the tuple up to *-isomorphism, since by [DDSS17, Proposition 5.5], $\operatorname{d}_{\mathrm{mr}}(A,B)=0$ if and only if there is unital completely isometric map sending $A_{i}$ to $B_{i}$ for $i=1,\ldots,d$ (thus the matrix range is a complete invariant of the operator system generated by a tuple, and it is known that completely isometric operator systems may generate non-isomorphic C*-algebras). In fact, [DDSS17, Proposition 5.5] gives more generally that $\mathcal{W}(B)\subseteq\mathcal{W}(A)$ if and only if there exists a unital completely positive (UCP) map $\phi$ such that $B_{i}=\phi(A_{i})$ for $i=1,\ldots,d$.
Another measure of difference between tuples is given by the dilation distance [GPSS21] defined as follows. First, we define the dilation constant $c(u,v)$. Given two unitary tuples $u,v$ and a real number $c>0$, we write $u\prec cv$ if there exist two Hilbert spaces $\mathcal{H}\subseteq\mathcal{K}$ and two operator tuples $U\in B(\mathcal{H})^{d}$ and $V\in B(\mathcal{K})^{d}$, such that $u\sim U$, $v\sim V$ and
@@
Various aspects of the dilation distance were studied in [GPSS21]. For the Haagerup-Rørdam distance
@@
in [GPSS21, Theorem 2.6] it was proved that
@@
where $K$ is a universal constant (for further developments see [GS23] and [GS24]).
@@ 1.2. Overview of this paper
For some time now it has been an open problem to determine the precise value of $C_{d}$ for all $d$. It is not hard to show that $C_{d}\leq d$ [DDSS17, HKMS19]. This was improved to $C_{d}\leq\max\{d,2\sqrt{d}\}$ in [PSS18] and to $C_{d}\leq\sqrt{2d}$ in [Pas19]. The lower bound $C_{d}\geq\sqrt{d}$ follows from the selfadjoint case obtained in [PSS18]. For general $d$ the best bounds are $\sqrt{d}\leq C_{d}\leq\sqrt{2d}$. However, for $d=2$ and $d=3$, it was shown in [GS20] and [GPSS21], respectively, that $C_{d}>\sqrt{d}$. Thus, the lower bound for $C_{d}$ is not tight. In this paper we provide evidence that the upper bound is not tight, either. We focus on the case $d=2$, in which the currently known best upper bound $C_{2}\leq\sqrt{2\cdot 2}$ coincides with the easy bound $C_{d}\leq d$.
@@
In [GS21], we studied matrix ranges of random matrix tuples. We showed that if $T^{(N)}=(T_{1}^{(N)},\ldots,T_{d}^{(N)})$ is a random matrix ensemble and $t=(t_{1},\ldots,t_{d})$ a $d$-tuple of operators such that
@@
almost surely (see [GS21, Theorem 3.1]). When combined with known results on strong convergence of matrix ensembles, this result implies convergence of matrix ranges of these matrix ensembles. For example, if $X^{(N)}$ are tuples of independent matrices from a Wigner ensemble, then for all $n$,
@@
More interesting for our purposes are ensembles of unitaries. Recall that a free tuple of Haar unitaries is a $d$-tuple $u_{\mathrm{f}}=(u_{\mathrm{f}1},\ldots,u_{\mathrm{f}d})$ of unitaries in a C*-probability space $(\mathcal{A},\tau)$ which are freely independent and satisfy $\tau(u_{\mathrm{f}i}^{k})=0$ for all $i$ and all $k\in\mathbb{Z}\setminus\{0\}$. By [GS21, Theorem 3.4], if $U^{(N)}=(U^{(N)}_{1},\ldots,U^{(N)}_{d})$ are $d$ independent $N\times N$ unitaries sampled from the Haar measure on $\mathcal{U}_{N}$, and $u_{\mathrm{f}}=(u_{\mathrm{f}1},\ldots,u_{\mathrm{f}d})$ is a free Haar unitary tuple, then the matrix range $\mathcal{W}(U^{(N)})$ converges almost surely levelwise in the Hausdorff metric to $\mathcal{W}(u_{\mathrm{f}})$, that is, for all $n$,
@@
On the other hand, by [GPSS21, Corollary 3.8], we have the exact value
@@
It is worth recalling that by Theorems 3.9 and 3.10 in [GPSS21], we have the estimate
@@
Plugging the upper bound in Equation˜5 one recovers the upper bound $C_{d}\leq\sqrt{2d}$, which was obtained in [Pas19]; in the case $d=2$ this is equal to the trivial bound $C_{2}=2$. On the other hand, in the case $d=2$, the lower bound gives $\sqrt{2}$ which is what the data suggests.
@@ 2. Levelwise convergence and matrix range distance
Then there are sequences $N_{k},M_{k}\xrightarrow{k\to\infty}\infty$ and $r_{k}<r$ such that $\xi^{(N_{k})}\prec r_{k}\xi^{(M_{k})}$ for all $k\in\mathbb{N}$ or, equivalently (by [DDSS17, Proposition 5.5]), $\mathcal{W}_{n}(\xi^{(N_{k})})\subset r_{k}\mathcal{W}_{n}(\xi^{(M_{k})})$ for all $n,k\in\mathbb{N}$. By passing to a subsequence, we may assume that $r_{k}\to r_{\infty}\leq r$. It follows from the assumption of levelwise convergence that $\mathcal{W}_{n}(\xi^{(\infty)})\subset r_{\infty}\mathcal{W}_{n}(\xi^{(\infty)})$ for all $n\in\mathbb{N}$, i.e. $\xi^{(\infty)}\prec r_{\infty}\eta^{(\infty)}$, which implies to $c(\xi^{(\infty)},\eta^{(\infty)})\leq r_{\infty}\leq r$. Therefore, the claimed inequality must hold. ∎
@@
The limit inferior in Theorem˜2.1 cannot be improved to a limit. Consider $\xi^{(N)}$ be a sequence of tuples such that $\mathcal{W}(\xi^{(N)})$ converges levelwise, but not uniformly to $\mathcal{W}(\xi^{(\infty)})$ (for the existence of examples of such tuples, see [PP21]), and let $\eta^{(\infty)}=\eta^{(N)}=\xi^{(\infty)}$. Then by passing to a subsequence, we find that
@@
showing that weak inequality cannot, in general, be replaced by an equality in the theorem. Similarly, one can show that the limit inferior cannot be replaced by a limit.
@@
By the strong asymptotic freeness property and [GS21, Theorem 3.1], $\mathcal{W}(T^{(N)})$ converges levelwise to $\mathcal{W}(u_{\mathrm{f}})$. Strong asymptotic freeness for the first two named cases is proved in [CM14]. Strong asymptotic freeness of permutation matrices in standard representation (i.e. acting on the orthogonal complement of the Perron-Frobenius Eigenvector) is shown in [BC19]. This is good enough, because $T=\pi_{\mathrm{std}}(T)\oplus\pi_{\mathrm{trivial}}(T)$ and the $\pi_{\mathrm{trivial}}(T_{i}^{(N)})=1$ commute. ∎
@@
Corollary˜2.4 can also be shown to hold without recourse to matrix ranges, using the fact that by the implication $(i)\Rightarrow(ii)$ in [GS21, Theorem 2.3], if Equation˜2 holds for all *-polynomials, then it also holds for all matrix valued *-polynomials. We omit the details.
@@
The uniform convergence of $\mathcal{W}^{k\text{-$\max$}}(X)$ and $\mathcal{W}^{k\text{-$\min$}}(X)$ to $\mathcal{W}(X)$ was characterized by Passer and Paulsen in operator system terms (lifting property and 1-exactness) in [PP21, Theorems 3.3 and 3.7].
@@ 3.2. Algorithm for computing $c(\mathsf{U},\mathsf{N})$
The following algorithm for computing the dilation constant $c(\mathsf{U},\mathsf{N})$ is an adaptation of the algorithm by Helton, Klep and McCullough [HKM13] for determining whether there exists a UCP map between two sets of matrices. In this form it appeared first in [Sha19].
@@
Note that the constraints ˜9 and 10 mean that the map $E_{jj}\mapsto\mathsf{C}_{j}$ (for $j=1,\ldots,k^{2}$) defines a UCP map from the diagonal $k^{2}\times k^{2}$ matrices to $M_{N}(\mathbb{C})$; by Arveson’s extension theorem, this map extends to a UCP map $\Phi\colon M_{k^{2}}(\mathbb{C})\to M_{N}(\mathbb{C})$. The constraint ˜11 means that $\Phi(\mathsf{N}_{i})=r\mathsf{U}_{i}$ for $i=1,2$. Thus, a solution to the above semidefinite program is the maximal $r$ such that there exists a UCP map $\Phi$ such that $\Phi(\mathsf{N})=r\mathsf{U}$. The dilation constant $c(\mathsf{U},\mathsf{N})$ is therefore given by $c(\mathsf{U},\mathsf{N})=r^{-1}$.

### v2

@@ 1. Introduction
In recent years, dilation theoretic techniques and the framework of matrix convex sets have found new and significant applications to quantum information [5, 6, 4], optimization and control [23], mathematical physics [2, 20], the theory of operator systems and operator algebras [10, 15, 32], and more [12, 13, 14, 25]. The goal of this paper is to advance our knowledge regarding an open problem in this area: determining the universal commuting dilation constant $C_{2}$, i.e. the smallest number $c$ such that every pair of contractions is the simultaneous compression of commuting normal operators of norm at most $c$. By combining rigorous limit theorems with numerical experimentation, we provide evidence that $C_{2}\leq 2\sqrt{\frac{2}{3}}<2$.
@@ 1.1. Preliminaries and background
The one sided matrix range distance $\operatorname{d}_{\mathrm{mr}}(B\to A)$ was discussed before in [11, Section 5] where it was denoted by $\delta_{\mathcal{W}(A)}(\mathcal{W}(B))$. The matrix range distance is then given by
@@
The matrix range distance is a metric only when considered on classes of certain rigid tuples, for example on the set $\mathcal{U}(d)$ of unitary $d$-tuples. For general operator tuples it does not determine the tuple up to *-isomorphism, since by [11, Proposition 5.5], $\operatorname{d}_{\mathrm{mr}}(A,B)=0$ if and only if there is unital completely isometric map sending $A_{i}$ to $B_{i}$ for $i=1,\ldots,d$ (thus the matrix range is a complete invariant of the operator system generated by a tuple, and it is known that completely isometric operator systems may generate non-isomorphic C*-algebras). In fact, [11, Proposition 5.5] gives more generally that $\mathcal{W}(B)\subseteq\mathcal{W}(A)$ if and only if there exists a unital completely positive (UCP) map $\phi$ such that $B_{i}=\phi(A_{i})$ for $i=1,\ldots,d$.
Another measure of difference between tuples is given by the dilation distance [17] defined as follows. First, we define the dilation constant $c(u,v)$. Given two unitary tuples $u,v$ and a real number $c>0$, we write $u\prec cv$ if there exist two Hilbert spaces $\mathcal{H}\subseteq\mathcal{K}$ and two operator tuples $U\in B(\mathcal{H})^{d}$ and $V\in B(\mathcal{K})^{d}$, such that $u\sim U$, $v\sim V$ and
@@
Various aspects of the dilation distance were studied in [17]. For the Haagerup-Rørdam distance
@@
in [17, Theorem 2.6] it was proved that
@@
where $K$ is a universal constant (for further developments see [20, 21, 16]).
@@ 1.2. Overview of this paper
For some time now it has been an open problem to determine the precise value of $C_{d}$ for all $d$. It is not hard to show that $C_{d}\leq d$ [11, 23]. This was improved to $C_{d}\leq\max\{d,2\sqrt{d}\}$ in [33] and to $C_{d}\leq\sqrt{2d}$ in [34]. The lower bound $C_{d}\geq\sqrt{d}$ follows from the selfadjoint case obtained in [33]. For general $d$ the best bounds are $\sqrt{d}\leq C_{d}\leq\sqrt{2d}$. However, for $d=2$ and $d=3$, it was shown in [18] and [17], respectively, that $C_{d}>\sqrt{d}$. Thus, the lower bound for $C_{d}$ is not tight. In this paper we provide evidence that the upper bound is not tight, either. We focus on the case $d=2$, in which the currently known best upper bound $C_{2}\leq\sqrt{2\cdot 2}$ coincides with the easy bound $C_{d}\leq d$.
@@
In [19], we studied matrix ranges of random matrix tuples. We showed that if $T^{(N)}=(T_{1}^{(N)},\ldots,T_{d}^{(N)})$ is a random matrix ensemble and $t=(t_{1},\ldots,t_{d})$ a $d$-tuple of operators such that
@@
almost surely (see [19, Theorem 3.1]). When combined with known results on strong convergence of matrix ensembles, this result implies convergence of matrix ranges of these matrix ensembles. For example, if $X^{(N)}$ are tuples of independent matrices from a Wigner ensemble, then for all $n$,
@@
More interesting for our purposes are ensembles of unitaries. Recall that a free tuple of Haar unitaries is a $d$-tuple $u_{\mathrm{f}}=(u_{\mathrm{f}1},\ldots,u_{\mathrm{f}d})$ of unitaries in a C*-probability space $(\mathcal{A},\tau)$ which are freely independent and satisfy $\tau(u_{\mathrm{f}i}^{k})=0$ for all $i$ and all $k\in\mathbb{Z}\setminus\{0\}$. By [19, Theorem 3.4], if $U^{(N)}=(U^{(N)}_{1},\ldots,U^{(N)}_{d})$ are $d$ independent $N\times N$ unitaries sampled from the Haar measure on $\mathcal{U}_{N}$, and $u_{\mathrm{f}}=(u_{\mathrm{f}1},\ldots,u_{\mathrm{f}d})$ is a free Haar unitary tuple, then the matrix range $\mathcal{W}(U^{(N)})$ converges almost surely levelwise in the Hausdorff metric to $\mathcal{W}(u_{\mathrm{f}})$, that is, for all $n$,
@@
On the other hand, by [17, Corollary 3.8], we have the exact value
@@
It is worth recalling that by Theorems 3.9 and 3.10 in [17], we have the estimate
@@
Plugging the upper bound in Equation˜5 one recovers the upper bound $C_{d}\leq\sqrt{2d}$, which was obtained in [34]; in the case $d=2$ this is equal to the trivial bound $C_{2}=2$. On the other hand, in the case $d=2$, the lower bound gives $\sqrt{2}$ which is what the data suggests.
@@ 2. Levelwise convergence and matrix range distance
Then there are sequences $N_{k},M_{k}\xrightarrow{k\to\infty}\infty$ and $r_{k}<r$ such that $\xi^{(N_{k})}\prec r_{k}\eta^{(M_{k})}$ for all $k\in\mathbb{N}$ or, equivalently (by [11, Proposition 5.5]), $\mathcal{W}_{n}(\xi^{(N_{k})})\subset r_{k}\mathcal{W}_{n}(\eta^{(M_{k})})$ for all $n,k\in\mathbb{N}$. By passing to a subsequence, we may assume that $r_{k}\to r_{\infty}\leq r$. It follows from the assumption of levelwise convergence that $\mathcal{W}_{n}(\xi^{(\infty)})\subset r_{\infty}\mathcal{W}_{n}(\eta^{(\infty)})$ for all $n\in\mathbb{N}$, i.e. $\xi^{(\infty)}\prec r_{\infty}\eta^{(\infty)}$, which implies to $c(\xi^{(\infty)},\eta^{(\infty)})\leq r_{\infty}\leq r$. Therefore, the claimed inequality must hold. ∎
@@
Consider a sequence $\xi^{(N)}$ of tuples such that $\mathcal{W}(\xi^{(N)})$ converges levelwise, but not uniformly to $\mathcal{W}(\xi^{(\infty)})$ (for the existence of examples of such tuples, see [32]), and let $\eta^{(\infty)}=\eta^{(N)}=\xi^{(\infty)}$. Then by passing to a subsequence, we find that
@@
showing that weak inequality cannot, in general, be replaced by an equality in Theorem˜2.1. Similarly, one can show that the limit inferior cannot be replaced by a limit; indeed, with $\xi^{(N)}$ as before, the double sequence $\mathrm{d_{mr}}(\xi^{(N)},\xi^{(M)})$ cannot converge.
@@
By the strong asymptotic freeness property and [19, Theorem 3.1], $\mathcal{W}(T^{(N)})$ converges levelwise to $\mathcal{W}(u_{\mathrm{f}})$, so Equation (7) follows from Theorem 2.1. Strong asymptotic freeness for the first two named cases is proved in [9]. Strong asymptotic freeness of permutation matrices in standard representation (i.e. acting on the orthogonal complement of the Perron-Frobenius Eigenvector) is shown in [7]. This is good enough, because $T=\pi_{\mathrm{std}}(T)\oplus\pi_{\mathrm{trivial}}(T)$ and the $\pi_{\mathrm{trivial}}(T_{i}^{(N)})=1$ commute. ∎
@@
Corollary˜2.4 can also be shown to hold without recourse to matrix ranges, using the fact that by the implication $(i)\Rightarrow(ii)$ in [19, Theorem 2.3], if Equation˜2 holds for all *-polynomials, then it also holds for all matrix valued *-polynomials. We omit the details.
@@
The uniform convergence of $\mathcal{W}^{k\text{-$\max$}}(X)$ and $\mathcal{W}^{k\text{-$\min$}}(X)$ to $\mathcal{W}(X)$ was characterized by Passer and Paulsen in operator system terms (lifting property and 1-exactness) in [32, Theorems 3.3 and 3.7].
We close this section with a lengthy remark on the selfadjoint case.
Remark 2.8.
Consider the dilation constants $\alpha_{d},\beta_{d},\gamma_{d},\delta_{d}$ defined as follows:
  1. (1)
    $\alpha_{d}$ is the smallest constant such that every $d$ selfadjoint contractions can be dilated to a $d$-tuple of commuting selfadjoint operators of norm at most $\alpha_{d}$.
  2. (2)
    $\beta_{d}$ is the smallest constant such that every $d$ selfadjoint contractions can be dilated to a $d$-tuple of commuting selfadjoint operators of the form $\beta_{d}\timesunitary$.
  3. (3)
    $\gamma_{d}$ is the smallest constant such that every $d$ selfadjoint contractions can be dilated to a $d$-tuple of commuting normal operators of norm at most $\gamma_{d}$.
  4. (4)
    $\delta_{d}$ is the smallest constant such that every $d$ selfadjoint contractions can be dilated to a $d$-tuple of commuting operators of the form $\delta_{d}\timesunitary$.
Tautologically, $\alpha_{d}\geq\gamma_{d}$, but since $A\prec B$ and $A=A^{*}$ imply that $A\prec\frac{1}{2}(B+B^{*})$, we have that $\alpha_{d}\leq\gamma_{d}$ thus $\alpha_{d}=\gamma_{d}$. Likewise, $\alpha_{d}\geq\beta_{d}$ and $\gamma_{d}\geq\delta_{d}$, but by “pushing the spectrum to the extreme" one can show that $\alpha_{d}=\beta_{d}$ and that $\gamma_{d}=\delta_{d}$ (see, e.g. [33, Proposition 2.3]); thus, all these constants are equal.
In a similar vein, one can show that if $x$ is a $d$-tuple of selfadjoint operators, then $c(x,u_{0}):=\min\{c:x\prec cu_{0}\}$ is equal to the minimal constant $c$ such that $x\prec cv$ for a tuple $v$ of commuting selfadjoint unitaries, and also equal to the minimal $c$ such that $x\prec cy$ for a commuting tuple $y$ of selfadjoint contractions.
By [33, Theorem 6.7] we know that $\alpha_{d}=\sqrt{d}$, thus, numerical experiments are not needed for shedding light on this constant. Moreover, the sharpness of $\sqrt{d}$ is exhibited already by a tuple of anti-commuting unitaries. In particular, for $d=2$, we have that the universal commuting selfadjoint dilation constant is achieved already as $c(F,u_{0})=\sqrt{2}$ for $F_{1}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$, $F_{2}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$. The referee asked whether the dilation constants for pairs of selfadjoint random $N\times N$ unitaries converge. If $D^{(N)}$ is a sequence of diagonal selfadjoint unitary matrices with empirical distribution converging to $\frac{1}{2}\delta_{-1}+\frac{1}{2}\delta_{1}$, and if $U^{(N)}_{1}$ and $U^{(N)}_{2}$ are two independent random unitary matrices, and if we define $R_{i}^{(N)}=U^{(N)}_{i}D^{(N)}(U^{(N)}_{i})^{*}$ for $i=1,2$, then it follows from [9, Theorem 1.4] that the pair of selfadjoint unitaries $R^{(N)}=(R_{1}^{(N)},R_{2}^{(N)})$ is asymptotically free and converges strongly to a pair of free Bernoulli variables $r=(r_{1},r_{2})$. Now [19, Theorem 3.1] shows that $\mathcal{W}(R^{(N)})\to\mathcal{W}(r)$ levelwise so that Theorem 2.1 applies and we find that
$\liminf_{N\to\infty}c(R^{(N)},u_{0})\geq c(r,u_{0}).$
To estimate $c(r,u_{0})$, note that $r_{1}$ and $r_{2}r_{1}r_{2}$ are freely independent Bernoulli variables, and therefore their sum $r_{1}-r_{2}r_{1}r_{2}$ is distributed according to the arcsine distribution on [2,2][-2,2] (see [29, Example 12.8]). It follows that the norm of the commutator $r_{1},r_{2}]=r_{1}r_{2}-r_{2}r_{1}$ is $\|[r_{1},r_{2}]\|=\|r_{1}-r_{2}r_{1}r_{2}\|=2$. Now the proof of [17, Theorem 3.9] (with $u=r_{1}$ and $v=r_{2}$) shows that $c(r,u_{0})\geq\sqrt{2}$. On the other hand, $c(R^{(N)},u_{0})\leq\sqrt{2}$ because $\alpha_{2}=\delta_{2}=\sqrt{2}$, and so $\limsup_{N\to\infty}c(R^{(N)},u_{0})\leq\sqrt{2}$. Putting the two inequalities together we conclude that
$\lim_{N\to\infty}c(R^{(N)},u_{0})=c(r,u_{0})=\sqrt{2},$
as the referee suspected.
Another natural question to ask about the selfadjoint case is the following. We know that if $X^{(N)}$ is a $d$-tuple of (appropriately normalized) independent random Wigner $N\times N$ matrices satisfying some additional assumptions (in particular, if they are sampled from the GUE ensemble), then $X^{(N)}$ converges strongly to the tuple $s=(s_{1},\ldots,s_{d})$ of free semicircular operators. We know from [19, Theorem 3.2] that $\mathcal{W}(X^{(N)})\to\mathcal{W}(s)$. Therefore, by Theorem 2.1, we have that
$c(s,u_{0})\leq\liminf_{N\to\infty}c(X^{(N)},u_{0}).$
However, we do not know much about $c(s,u_{0})$. Rudimentary numerical experiments that we carried out (for values $N=10,50,100,500,1000$) suggest that $\lim_{N\to\infty}c(X^{(N)}/2,u_{0})=1$ and combining with Theorem 2.1, this will imply that $c(s/2,u_{0})=1$ because it cannot be smaller than $1$. This would suggest the plausibility of the containment $\mathcal{W}(s/2)\subsetneq\mathcal{W}^{\operatorname{min}}([-1,1]^{d})$ (compare with [19, Theorem 5.4], where it is established that $\mathcal{W}^{\operatorname{min}}(\overline{\mathbb{B}}_{d})\subsetneq\mathcal{W}(s/2)$, where $\overline{\mathbb{B}}_{d}$ is the Euclidean unit ball in $\mathbb{R}^{d}$). It would be interesting to prove this rigorously.
@@ 3.2. Algorithm for computing $c(\mathsf{U},\mathsf{N})$
The following algorithm for computing the dilation constant $c(\mathsf{U},\mathsf{N})$ is an adaptation of the algorithm by Helton, Klep and McCullough [24] for determining whether there exists a UCP map between two sets of matrices. In this form it appeared first in [38].
@@
Note that the constraints ˜10 and 11 mean that the map $E_{jj}\mapsto\mathsf{C}_{j}$ (for $j=1,\ldots,k^{2}$) defines a UCP map from the diagonal $k^{2}\times k^{2}$ matrices to $M_{N}(\mathbb{C})$; by Arveson’s extension theorem, this map extends to a UCP map $\Phi\colon M_{k^{2}}(\mathbb{C})\to M_{N}(\mathbb{C})$. The constraint ˜12 means that $\Phi(\mathsf{N}_{i})=r\mathsf{U}_{i}$ for $i=1,2$. Thus, a solution to the above semidefinite program is the maximal $r$ such that there exists a UCP map $\Phi$ such that $\Phi(\mathsf{N})=r\mathsf{U}$. The dilation constant $c(\mathsf{U},\mathsf{N})$ is therefore given by $c(\mathsf{U},\mathsf{N})=r^{-1}$.